OpenStudy (anonymous):

Four of the angles of a pentagon measure 85°, 110°, 135°, and 95°. Find the measure of the missing angle. A) 115° B) 95° C) 135° D) 85°

OpenStudy (hba):

\[\huge \color{red}{\text{Welcome \to/ Open Study}}\ddot\smile\]

OpenStudy (hba):

How many sides does a pentagon have @HaileyroX! ?

OpenStudy (hba):

Please respond @HaileyroX! ?

OpenStudy (anonymous):

Five

OpenStudy (hba):

Well, Sum of interior angle=(n-2)*180 Here, n=5 So, Sum of interior angles=(5-2)*180 Now calculate :)

OpenStudy (anonymous):

is it C

OpenStudy (hba):

Did you calculate that ?

OpenStudy (anonymous):

ya

OpenStudy (hba):

What did you get ?

OpenStudy (hba):

@HaileyroX! ?